Find the number of possible 5-card hands that contain the cards specified. The cards are taken from a standard 52 dard deck.
A. 3 face cards and 2 cards that are not face cards. I have no idea what a face card is. I don't really play cards that often and know the terms.

Answer by shree840(260)

there are 52 cards jack queen king and aces are face cards
so there are 4*4=16 face cards
there are 52-16 =36 non face card
prob of face card 16/52 * 15/51*14/50
prob of two non face card =36/49 * 35/48
total prob =)16/52)*(15/51)*(14/50)*(36/49)*(35/48)